OpenStudy (anonymous):

Solve the system of equations by graphing. 3x-y=3, 4x+y=4

OpenStudy (jim766):

solve each equation for y 3x - y = 3 -3x -3x -y = -3x + 3 --- --- --- -1 -1 -1 y = -3x -3 Can you do the other one?

OpenStudy (anonymous):

i just need two plot points

OpenStudy (anonymous):

to graph them

OpenStudy (jdoe0001):

or use a graphing tool, and check where both "lines" meet/intersect, that'd be the value of Y for the solution, the common Y

OpenStudy (anonymous):

Refer to the attachment.